# Training-Free Generative Modeling via Kernelized Stochastic Interpolants

### v1

@@ 1 Introduction
Generative diffusion models based on dynamical transport of measure—including score-based models (Song et al., 2020), flow matching (Lipman et al., 2022), rectified flows (Liu et al., 2022), and stochastic interpolants (Albergo & Vanden-Eijnden, 2022; Albergo et al., 2023)—require learning a time-dependent drift or score function, typically by training a neural network on a regression objective. This training step is computationally expensive.
@@
Our approach builds on stochastic interpolants (Albergo & Vanden-Eijnden, 2022; Albergo et al., 2023), reformulating them as kernel methods (Muandet et al., 2017). The optimal diffusion coefficient $D_{t}^{*}$ we use was derived in Ma et al. (2024); Chen et al. (2024); a similar expression appears in the stochastic optimal control setting (Havens et al., 2025). Closely related is Moment-Guided Diffusion (MGD) (Lempereur et al., 2025), which also uses feature maps within stochastic interpolants but imposes moment constraints $\mathbb{E}[\phi(X_{t})]=\mathbb{E}[\phi(I_{t})]$ to sample maximum entropy distributions. MGD requires $D_{t}\to\infty$ to reach the maximum entropy limit, whereas our approach uses the finite optimal $D_{t}^{*}$ to minimize generation error. Our method is simpler—solving a linear regression beforehand rather than enforcing constraints throughout the dynamics—but offers less control: MGD allows explicit specification of which statistics to match, while our method implicitly targets the full data distribution through drift regression.
RKHS formulations of diffusion models appear in Maurais & Marzouk (2024); Yi et al. (2024), using standard kernels (Gaussian, Laplacian) rather than problem-specific feature maps. MMD-based generative methods (Gretton et al., 2012; Galashov et al., 2024; Bortoli et al., 2025) require engineering feature maps; we use $\nabla\phi$ directly from existing models or transforms.
@@ 2.5 Choice of Feature Map
For time series ($d=T$) or random fields ($d=T_{1}\times T_{2}$), the wavelet scattering transform (Mallat, 2012) provides a natural feature map. The scattering coefficients $S[p]x$ at various paths $p$ (encoding scales, orientations) give a finite-dimensional, Lipschitz-continuous, translation-invariant representation. Setting $\phi_{p}(x)=S[p]x$ and computing $\nabla\phi_{p}(x)=\nabla_{x}S[p]x$ via backpropagation provides the feature gradients. This choice is well-suited to processes with multiscale structure, such as textures and turbulence (Bruna & Mallat, 2019).
@@ Acknowledgments
This work was supported by PR[AI]RIE-PSAI-ANR-23-IACL-0008 and the DRUIDS projet ANR-24-EXMA-0002. It was granted access to the HPC resources of IDRIS under the allocations 2025-AD011016159R1 and 2025-A0181016159 made by GENCI.

### v2

@@ 1 Introduction
Generative diffusion models based on dynamical transport of measure—including score-based models (Song et al., 2020), flow matching (Lipman et al., 2022), rectified flows (Liu et al., 2022), and stochastic interpolants (Albergo and Vanden-Eijnden, 2022; Albergo et al., 2023)—require learning a time-dependent drift or score function, typically by training a neural network on a regression objective. This training step is computationally expensive.
@@
Our approach builds on stochastic interpolants (Albergo and Vanden-Eijnden, 2022; Albergo et al., 2023), reformulating them as kernel methods (Muandet et al., 2017). The optimal diffusion coefficient $D_{t}^{*}$ we use was derived in Ma et al. (2024); Chen et al. (2024); a similar expression appears in the stochastic optimal control setting (Havens et al., 2025). Closely related is Moment-Guided Diffusion (MGD) (Lempereur et al., 2025), which also uses feature maps within stochastic interpolants but imposes moment constraints $\mathbb{E}[\phi(X_{t})]=\mathbb{E}[\phi(I_{t})]$ to sample maximum entropy distributions. MGD requires $D_{t}\to\infty$ to reach the maximum entropy limit, whereas our approach uses the finite optimal $D_{t}^{*}$ to minimize generation error. Our method is simpler—solving a linear regression beforehand rather than enforcing constraints throughout the dynamics—but offers less control: MGD allows explicit specification of which statistics to match, while our method implicitly targets the full data distribution through drift regression.
RKHS formulations of diffusion models appear in Maurais and Marzouk (2024); Yi et al. (2024), using standard kernels (Gaussian, Laplacian) rather than problem-specific feature maps. MMD-based generative methods (Gretton et al., 2012; Galashov et al., 2024; Bortoli et al., 2025) require engineering feature maps; we use $\nabla\phi$ directly from existing models or transforms.
@@ 2.5 Choice of Feature Map
For time series ($d=T$) or random fields ($d=T_{1}\times T_{2}$), the wavelet scattering transform (Mallat, 2012) provides a natural feature map. The scattering coefficients $S[p]x$ at various paths $p$ (encoding scales, orientations) give a finite-dimensional, Lipschitz-continuous, translation-invariant representation. Setting $\phi_{p}(x)=S[p]x$ and computing $\nabla\phi_{p}(x)=\nabla_{x}S[p]x$ via backpropagation provides the feature gradients. This choice is well-suited to processes with multiscale structure, such as textures and turbulence (Bruna and Mallat, 2019).
@@ Acknowledgments
This work was supported by PR[AI]RIE-PSAI-ANR-23-IACL-0008 and the DRUIDS projet ANR-24-EXMA-0002.